SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Nov.1, 2005)

Attendees: Kris Chaisanguanthum, Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dong Su, Dan Walker.

Operations

  • Still no data taking yet. The DCH chamber is scheduled to have gas fill in tomorrow, so we are expecting to be able to test the new version of the TSF firmware at IR2 after the DCH electronics checked out okay.

EMT

  • The EMT tests of the prescale factor reading from the configuration database at IR2 went well last Thursday. Some COSMICS_NODCH runs were taken with prescale factors of 1, 2, 3, 500, and the xtc file for these runs all checked out fine, with the correct number of prescaled events. Currently, Dan leaves the prescale factor 3 in the configuration. Su Dong agrees that this is a good number to use for now. The next step is to wait for physics runs to check the impact of different prescale factors.
  • We noticed that in the cosmics run with a prescale of 500 last week, the fastmon has very low statistics. Su Dong suggested that we could do monitoring with EMT tower in L3 if we need more statistics there.
  • Dan is working on OPR monitoring to flag digi to indicate if the event is prescaled. In simulation, there should be no prescale.

DCT/DCZ

  • We were able to use the test stand to take DCH-DCT calibrations successfully with the latest version of the TSF firmware. This version passes all tests done at the test stand. We are now waiting for tests at IR2.
  • Fang Fang would like to load an updated version of the DCZ code for production soon (following up from last week's minutes). Selina sugggested this to be tested at the test stand first.

L3

  • Al found the odd setup with the L3 monitoring plots for JAS which makes it hard for someone who is not so familiar with the setup to make any changes at all. Selina will talk to Prafulla about this.

Data Quality and Monitoring

  • Doing Run 5 QA now. Nothing much to report.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.